How to you solve this? Solve the equation x(m+2)+y(m-8)=z(m-4),for m, assuming that x+y-z does not = 0. I would be very greatful if you could explain it to me!
Answer by DrBeeee(684)

Firstly move the z-term to the left side of the equation then multiply out the parenthetical expressions, yielding
(1) mx + 2x + my - 8y - mz + 4z = 0
Now factor m out of term containing an m and get
(2) m*(x + y - z) + 2x - 8y + 4z = 0
Now factor a two out of the last three terms, and move the result to the right hand side and get
(3) m*(x+y-z) = -2*(x-4y+2z)
Now divide both sides by the coefficient of m on the left side and yield the answer
(4) m = -2/[x+y-z] ; x+y-z not equal to zero